• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Mohamed Hafez

    Log entry time 00:46:02 on March 03, 2012

    Entry number 362812

    keyword=Compton cavity power restored

    The Compton cavity laser power has been restored after changing the fiber amplifier. The cavity lock at 9 kWatt.



    A copy of this log entry has been emailed to: nanda@jlab.org, mhafez@jlab.org



    Figure 1

    Cavity power 9 kW stability